OpenStudy (anonymous):

Given two rectangles, one with length = l and width = w, and the second rectangle with length = 2l and width = 4w. How much larger will the perimeter of the second rectangle be than the perimeter of the first rectangle? a)p2-p1=2*(1=3w) b)P2 = 2P1 + w c)P2 = 2P1 + 4w d)P2 = P1 + w please help !

OpenStudy (shamil98):

The perimeter of a rectangle is given by: P = 2L + 2W Let's calculate the first rectangle's perimeter.. it would just be P_1 = 2L + 2W Now for the second one: P_2 = 2(2L) + 2(4w) P_ 2 = 4L + 8w

OpenStudy (shamil98):

I think option C would best full fill this.. \[\Huge P_2 = 2P_1 + 4W\] \[\Huge P_2 = 2(2L + 2W) + 4w\] \[\Huge P_2 = 4L + 4W + 4W\] \[\Huge P_2 = 4L + 8W\]